Word problem to solve; thanks
purchased a combination of 25 air condition and heating systems for $440,000.
One single AC system cost 22,000 and single heat system cost 11,000. How many of each did they buy
Steps
1. assign variables
2. write algebraic model
3. solve model
4... answer
Thanks
bb

Answer by stanbon(75887)

purchased a combination of 25 air condition and heating systems for $440,000.
One single AC system cost 22,000 and single heat system cost 11,000. How many of each did they buy
-----------
Let # of AC units purchased be "x"; Value of these is 22000x dollars
----------
# of heat units purchased is "25-x"; Value of these is 11000(25-x)=275000-11000x
------------------
EQUATION:
value + value = 440,00 dollars
22000x + 275000 - 11000x = 440000
11000x = 165000
x = 15 (# of AC units purchased)
25-x = 10 (# of heating units purchased)
=============
Cheers,
Stan H.